4.A spring mass harmonic oscillator consists of a 0.2kg mass sphere connected vertically with a spring of negligible mass and force constant of 6kN / m. The spring is released from rest 3cm from the equilibrium position. Calculate:
(a) The energy of the spring,
(b) The potential energy a when the compression of the spring is 1/3 of the amplitude,
(c) Kinetic energy at this time.
Answer #1

4. m = 0.2 kg

k = 6000 N/m


(a) energy = k A^2 / 2
= (6000)(0.03^2)/2

= 2.7 J


(B) PE = k x^2 / 2 = (6000)(0.03/3)^2 / 2

= 0.3 J


(C) total energy = PE + KE

2.7 = 0.3 + KE

KE = 2.4 J